The sum of the squares of three consecutive ODD integers is 875. Find the integers
Answer by checkley77(12844)

LET X, (X+2), & (X+4) BE THE 3 CONSECUTIVE ODD INTEGERS.
X^2+(X+2)^2+(X+4)^2=875
X^2+X^2+4X+4+X^2+8X+16=875
3X^2+12X+4+16-875=0
3X^2+12X-855=0
3(X^2+4X-285)=0
3(X-15)(X+19)=0
X-15=0
X=15 ANSWER FOR THE FIRST INTEGER.
15=2=17 FOR THE SECOND.
15+4=19 FOR THE THIRD.
PROOF:
15^2+17^2+19^2=875
225+289+361=875
875=875